OpenStudy (anonymous):

One solution of the equation 2x^2 − x − 10 = 0 is 5/2 . What is the other solution?

OpenStudy (anonymous):

@Mertsj here it is!

OpenStudy (mertsj):

If 5/2 is an answer that means x=5/2

OpenStudy (mertsj):

Multiply both sides by 2 and you get 2x=5

OpenStudy (mertsj):

Subtract 5 from both sides and you get: 2x-5=0

OpenStudy (mertsj):

That tells you that one factor of your equation is 2x-5

OpenStudy (mertsj):